OpenStudy (anonymous):

1. y = -2/3x -2 2. –2x + 3y = -6 explain how the graphs of the two equations in your group are the same and how they are different.

OpenStudy (jackellyn):

slope intercept form is y=mx + b m is slope b is y intercept solve for y to get that form -2x + 3y = -6 3y = -6 +2x y = (-6+2x)/3 y = -2 + 2/3 x or y= 2/3x -2 use the info i gave you above to identify the slope and y intercept.

OpenStudy (anonymous):

thanks

OpenStudy (jackellyn):

no problem :)